If i buy a 5 kilogram packet of rice and i use 1500g of it. What fraction of the original amount of rice do i have left?
1
Expert's answer
2019-05-13T09:28:47-0400

(5000-1500)/5000=3500/5000=0.7=70%.

70% of the original amount left.